OpenStudy (anonymous):

the lim -3/x as x goes to 0 is the anserw DNE

OpenStudy (anonymous):

what is DNE

OpenStudy (anonymous):

dose not exist !

OpenStudy (anonymous):

is infinite

OpenStudy (anonymous):

\[\infty \]

OpenStudy (anonymous):

think of to this way, you are deviding 3 by smaller and smaller number, so get get more and more parts

jimthompson5910 (jim_thompson5910):

Since \[\Large \lim_{x\to0^{-}}\left(-\frac{3}{x}\right) = \infty\] and \[\Large \lim_{x\to0^{+}}\left(-\frac{3}{x}\right) = -\infty\] The limit as x approaches zero does not exist

OpenStudy (anonymous):